OpenStudy (anonymous):

Find all the values of r such that y=Ax^r satisfies the equation: y''+(x^-1)y'-y^3=0 Give a reason why other solutions must exist for this equation.

OpenStudy (anonymous):

y=x^r y'=rx^r-1 y''=r(r-1)x^r-2 r(r-1)x^r-2+x^-1rx^-1-(y^3)=0 r(r-1)x^-2+rx^-2-y^3=0 do you know how to get y^3 in terms or r?

OpenStudy (anonymous):

r^2-r+r=y^3 r^2=x^5 because divide everything by x^-2 and im taking it as A is a constant r^2=Ax^5 2ln(r)=5ln(x) ln(r)=5/2ln(x) r=e^x^5/2 but i dont know i never did this one in linear algebra/D.E's